OpenStudy (ggg):

if x=6x+3, 1/x=? solve

OpenStudy (anonymous):

u kidding?

OpenStudy (anonymous):

5x=-3 x=-3/5 1/x=-5/3(final answer) medal please

OpenStudy (ggg):

it is wrong